OpenStudy (anonymous):

A line perpendicular to y=1/4x+3 passes through the point (0,6). Which other point lies on the line? A) (4,-2) B) (3,-6) C) (0, -6) D) (-2, 2)

OpenStudy (anonymous):

can u tell me which chapter this ques is related to

OpenStudy (freckles):

So we need to find the line perpendicular to y=1/4 * x+3 that goes through (0,6) First job is to identify the slope of y=1/4 * x+3.

OpenStudy (anonymous):

i have no idea... i got a math packet for spring break and all it says is chapters 1 through 9

OpenStudy (anonymous):

ok...

OpenStudy (freckles):

Hint: The slope of y=mx+b is m. So the slope of y=1/4 * x+3 is?

OpenStudy (anonymous):

1/4?

OpenStudy (freckles):

Yes. Now to find the slope of the perpendicular line we must flip that fraction and then take the opposite of value of that flipping.

OpenStudy (freckles):

So when you flip 1/4 you get?

OpenStudy (anonymous):

4

OpenStudy (freckles):

and you take the opposite of that number you get?

OpenStudy (anonymous):

1/4?

OpenStudy (freckles):

flip is reciprocal opposite you change signs

OpenStudy (freckles):

1/4 is not the opposite of 4

OpenStudy (freckles):

-4 is the opposite of 4

OpenStudy (anonymous):

ooooo ok so its -1/4?

OpenStudy (freckles):

-4 is the slope of the perpendicular line.

OpenStudy (anonymous):

oooo ok

OpenStudy (freckles):

no

OpenStudy (freckles):

you flip 1/4 and then take the opposite of that result. the flipping of 1/4 is 4 the opposite of that flipping is -4.

OpenStudy (anonymous):

ok

OpenStudy (freckles):

So we know know the line is y=mx+b where m is -4 y=-4x+b

OpenStudy (freckles):

we need to find b

OpenStudy (freckles):

we know this line goes through (0,6)

OpenStudy (anonymous):

ok

OpenStudy (freckles):

(0,6) is actually the y-intercept

OpenStudy (freckles):

since the x value is 0 there

OpenStudy (freckles):

y=mx+b m is the slope b is the y-intercept

OpenStudy (freckles):

your b is ? take a guess

OpenStudy (anonymous):

6

OpenStudy (freckles):

yes so your line that is perpendicular to y=1/4*x+3 is y=-4*x+6

OpenStudy (anonymous):

ok

OpenStudy (freckles):

and that goes through (0,6) of course

OpenStudy (freckles):

now of the points listed which point satisfies y=-4x+6

OpenStudy (anonymous):

so can i basically put in the answer choices and see if it works?

OpenStudy (freckles):

For example, you are trying to see which point gives you the same thing on both sides. Will (4,-2) work? Replace x with 4 and y with -2. y=-4x+6 -2=-4(4)+6 -2=-16+6 -2=-10 so the point (4,-2) is definitely not on y=-4x+6 since we know -2=-10 is definitely not true. You are looking for a point that will make the equation true. That point that makes the equation true is a point on the line.

OpenStudy (anonymous):

Its B right?

OpenStudy (anonymous):

(3, -6)?

OpenStudy (freckles):

Yes actually. y=-4x+6 replace x with 3 and y with -6 -6=-4(3)+6 -6=-12+6 -6=-6 totally true so (3,-6) is totally on the line

OpenStudy (anonymous):

ok great thank you very much!

OpenStudy (freckles):

np